# Bias-Corrected Data Synthesis for Imbalanced Learning

### v1

@@ 3.1 Bias Correction for Risk Functions
By Assumption 1 (A1), $T$ represents the transformation from the majority distribution to the minority distribution: if $\bm{X}\sim\mathcal{P}_{0}$, then $T(\bm{X})\sim\mathcal{P}_{1}$. For synthetic distributions, the pair $(\tilde{\mathcal{P}}_{0},\tilde{\mathcal{P}}_{1})$ approximately preserves the same relation: $\tilde{\mathcal{P}}_{1}$ is close in $\mathcal{W}_{1}$ distance to $(\tilde{\mathcal{P}}_{0})_{\#T}$ up to $\varepsilon_{T}$. This ensures that the synthetic generator retains the transformation structure. (A2) ensures the Lipschitz continuity of the loss and the transformed loss. Small covariate perturbations change the loss after transformation $T$ by at most $L_{T}$ times the perturbation size. The stability guarantees that small input fluctuations do not lead to extremely large changes in loss. (A3) further bounds the discrepancy between the minority loss and the majority loss. This condition requires that the loss function $\ell_{f}$ is roughly symmetric in expectation under the majority distribution $\mathcal{P}_{0}$ and the minority distribution $\mathcal{P}_{1}$ up to an error bound $\varepsilon_{h}$. This property also extends to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ and its transformation $(\tilde{\mathcal{P}}_{0})_{\#T}$. It ensures that the transformation does not introduce excessive bias into loss evaluation across the two distributions. Under these assumptions, we now establish a high-probability upper bound for $|\hat{\Delta}_{1}-\hat{\Delta}_{0}|$.
Proposition 1.
@@
Suppose Assumption 1 holds and there exists $0<c_{1}<c_{2}<1/2$ such that $c_{1}\leq n_{1}/n\leq c_{2}$. Then for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
@@
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}+\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
@@
Specifically, when $n_{1}^{*}=O(n_{0}-n_{1})$, $\tilde{n}_{1}=n_{1}^{*}=O(n_{0}-n_{1})$, $n_{0c}=n_{1}^{*}=O(n_{0}-n_{1})$ and $\tilde{n}_{0}=n_{1}^{*}=O(n_{0}-n_{1})$, the result can be written as follows: For any $\alpha>0$ and some constant $C>0$, with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}+C\sqrt{\frac{\log(8/\alpha)}{n_{0}-n_{1}}}.$
@@
Proposition 1 provides an upper bound for the difference between the majority and minority groups. This upper bound does not directly depend on how accurately the synthetic distribution $\tilde{\mathcal{P}}_{0}$ recovers the true distribution $\mathcal{P}_{0}$. Instead, the discrepancy is controlled through three components: the loss gap in expectation $\varepsilon_{h}$, the transformation approximation error $\varepsilon_{T}$, and the perturbation fluctuation in the order of $1/\sqrt{n}$. This implies that even when the synthetic generator produces samples that poorly approximate the true minority distribution, the bias correction step is nevertheless able to keep the additional error within a well-defined bound. In practice, this means that the procedure remains stable and effective even for poorly performed synthetic generators, ensuring the reliability of the bias-corrected estimator.
@@
    The loss $\ell_{f}(\bm{x},y)$ is Lipschitz in $\bm{x}$ with constant $L_{\ell}$ uniformly over $y\in\{0,1\}$, i.e., $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|$ for any $\bm{x}_{1},\bm{x}_{2}\in\mathbb{R}^{d}$ and $y\in\{0,1\}$.
@@
    There exists a constant $C_{3}>0$ such that for any $\bm{x}_{1}\in{\rm supp}(\mathcal{P}_{1})$, any of its $K$ nearest neighbors $\bm{x}_{2}$, and any $u\in[0,1]$,
    $|\ell_{f}(\bm{x}_{1}+u(\bm{x}_{2}-\bm{x}_{1}),1)-\ell_{f}(\bm{x}_{1},1)|\geq C_{3}u\|\bm{x}_{2}-\bm{x}_{1}\|_{2}.$
Assumption 2 (A1) guarantees that the loss function $\ell_{f}$ is Lipschitz continuous with respect to $\bm{x}$ uniformly over $y\in\{0,1\}$. Assumption 2 (A2) ensures that the minority distribution has bounded support and a bounded density function on its support. Assumption 2 (A3) holds, for example, if $\ell_{f}(\cdot,1)$ is differentiable on ${\rm supp}(\mathcal{P}_{1})$ and its gradient satisfies $(\bm{x}_{2}-\bm{x}_{1})^{\top}\nabla\ell_{f}(\bm{z},1)\geq C_{3}\|\bm{x}_{2}-\bm{x}_{1}\|_{2}$ for any $\bm{z}$ on the segment between $\bm{x}_{1}$ and $\bm{x}_{2}$.
The following theorem shows that when the minority synthetic samples are generated by SMOTE, the induced bias cannot vanish too quickly. Specifically, there is a lower bound that scales with $(K/n_{1})^{1/d}$, reflecting the discrepancy between $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$.
Proposition 2.
Suppose Assumption 2 holds, and the synthetic minority samples are i.i.d. generated from $\tilde{\mathcal{P}}_{1}$ via SMOTE with parameter $K$. Then there exists a constant $c_{1}>0$, depending on $(d,C_{1},C_{2},C_{3},R)$, such that for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}|\geq c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}-\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Proposition 2 demonstrates that the bias introduced into the loss function by the synthetic samples is statistically non-negligible. This bias prevents the synthetic augmented loss function (2.2) from serving as a close substitution for the balanced loss function (2.5). This discrepancy between these loss functions leads to a noticeable difference between their corresponding minimizers, which will potentially reduce the performance of the trained classifier.
@@
where $\mathcal{F}=\{f:\mathcal{X}\to(0,1)\}$ represents the prediction function class. In practice, however, one does not have access to the population balanced loss $L^{*}$, but instead minimizes an empirical loss. With synthetic samples generated and the bias correction procedure, we have the empirical minimizer of the bias-corrected loss function (2.8):
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$
Next, we investigate the following question: how close is the empirical bias-corrected minimizer $\hat{f}^{\rm bc}$ to the population balanced minimizer $f^{*}$?
To answer this question, we consider a uniform assumption, which requires that the gap between the bias terms for the minority and majority groups remains controlled. Suppose that Assumption 1 is satisfied for all prediction functions $f\in\mathcal{F}$. The basic idea indicates that the bias observed in the minority group can be transferred to the majority group up to a controlled error $\varepsilon_{\rm BT}$, where
$\varepsilon_{\rm BT}=2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}.$
Intuitively, if $\varepsilon_{\rm BT}$ is small, then correcting for the minority bias using information from the majority group is reliable.
@@
Suppose Assumption 1 is satisfied for all $f\in\mathcal{F}$ and $\tilde{n}_{1}/(n_{0}-n_{1})\to 1$. Then for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$L^{*}(\hat{f}^{\rm bc})-L^{*}(f^{*})\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}$
$+\sqrt{\frac{\log(10/\alpha)}{2}}\left\{\frac{\pi_{1}/\pi_{0}}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\left(\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}\right)\right\}.$
@@ 3.2 Bias Correction for Multi-Task Learning
The following theorem establishes nonasymptotic bounds for three types of estimations: the raw MLE $\hat{\bm{\beta}}_{k}$, the synthetic augmented estimator $\tilde{\bm{\beta}}_{k}$, and the bias-corrected estimator $\tilde{\bm{\beta}}_{k}^{\rm bc}$. It also quantifies how these parameter errors propagate to the estimation of the shared low-rank structure.
@@
Under Assumption 3, for any $\alpha\in(0,1)$, there exist constants $C_{1},C_{2},C_{3}>0$ such that, with probability at least $1-\alpha$, the following properties hold simultaneously for each source $k$:
${\rm(i)}\quad$ $C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\max}(\bm{H}_{k})}\right)\leq\|\hat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}\leq C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\min}(\bm{H}_{k})}\right),$
${\rm(ii)}\quad$ $\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
$\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
${\rm(iii)}\quad$ $\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling},k}\right),$
where
$\varepsilon_{{\rm sampling},k}=$ $C_{3}R\sqrt{\log(10d/\alpha)}\left\{\frac{\pi_{1}/(2\pi_{0})}{\sqrt{n_{k1}}}+\frac{1}{2\sqrt{n_{k0}}}+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right)\right\}.$
Furthermore, consider the $d\times K$ true and estimated matrices as $\bm{M}=[\bm{\beta}_{1},\cdots,\bm{\beta}_{K}]$, $\hat{\bm{M}}^{\rm raw}=[\hat{\bm{\beta}}_{1}^{\rm raw},\cdots,\hat{\bm{\beta}}_{K}^{\rm raw}]$, $\hat{\bm{M}}^{\rm syn}=[\hat{\bm{\beta}}_{1}^{\rm syn},\cdots,\hat{\bm{\beta}}_{K}^{\rm syn}]$ and $\hat{\bm{M}}^{\rm bc}=[\hat{\bm{\beta}}_{1}^{(\rm bc)},\cdots,\hat{\bm{\beta}}_{K}^{(\rm bc)}]$. Let $\bm{U}$ be the matrix of leading $r$ left singular vectors of $\bm{M}$ and $D=\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})$ the spectral gap. Then, with the same probability,
${\rm(iv)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm raw},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(v)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm syn},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(vi)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm bc},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2}.$
@@
    The raw MLE $\hat{\bm{\beta}}_{k}^{\rm raw}$ converges at the standard parametric rate $n_{k}^{-1/2}$.
@@
    The synthetic estimator $\hat{\bm{\beta}}^{\rm syn}$ inherits an additional bias, which reflects the distributional mismatch between $\mathcal{P}_{k,0}$ and $\tilde{\mathcal{P}}_{k,0}$.
@@
    The bias-corrected estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$ removes the bias at the cost of extra sampling fluctuations and the residual class-difference bias $\varepsilon_{\rm BT}$.
@@
the bias correction parameter estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$ has smaller errors than the synthetic augmented parameter estimator $\hat{\bm{\beta}}^{\rm syn}$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed. In addition, when $\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})=O_{p}(n_{k}^{-1/2})$, the lower bound for the synthetic-augmented estimator error $\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|$ in Theorem 3.2 becomes negative and thus can be replaced by zero. Theorem 3.2 shows that the bias correction performs well, especially for “bad” synthetic generators.

### v2

@@ 3.1 Bias Correction for Risk Functions
By Assumption 3.4 (A1), the minority distribution $\mathcal{P}_{1}$ is approximately characterized as a transformation of the majority distribution through a mapping $T$, in the sense that the pushforward measure $(\mathcal{P}_{0})_{\#T}$ is close to $\mathcal{P}_{1}$ in $\mathcal{W}_{1}$ distance up to $\varepsilon_{T}$. Assumption 3.4 (A1) is imposed as a structural alignment condition rather than a requirement for an explicitly known or computable transformation. The transformation $T$ is introduced solely for theoretical analysis to characterize the degree of distributional similarity between majority and minority populations. In practice, our method does not require estimating or specifying $T$, and the validity of the results depends only on the existence of such a transformation with small Wasserstein discrepancy. For the synthetic distributions, the same transformation structure is approximately preserved: $\tilde{\mathcal{P}}_{1}$ remains close in $\mathcal{W}_{1}$ distance to $(\tilde{\mathcal{P}}_{0})_{\#T}$, with the discrepancy controlled by the same tolerance $\varepsilon_{T}$. This ensures that the synthetic generator retains the transformation structure between the majority and minority distributions. (A2) ensures the Lipschitz continuity of the loss and the transformed loss. Small covariate perturbations change the loss after transformation $T$ by at most $L_{\ell}$ times the perturbation size. The stability guarantees that small input fluctuations do not lead to extremely large changes in loss. (A3) further bounds the discrepancy between the minority loss and the majority loss. This condition requires that the loss function $\ell_{f}$ is roughly symmetric in expectation under the majority distribution $\mathcal{P}_{0}$ and the minority distribution $\mathcal{P}_{1}$ up to an error bound $\varepsilon_{h}$. This property also extends to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ and its transformation $(\tilde{\mathcal{P}}_{0})_{\#T}$. It ensures that the transformation does not introduce excessive bias into loss evaluation across the two distributions. Under these assumptions, we now establish a high-probability upper bound for $|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|$. Denote $\varepsilon_{\rm BT}=2L_{\ell}\varepsilon_{T}+2\varepsilon_{h}$ as the bias-transfer error.
Proposition 3.1.
@@
Suppose Assumptions 3.1-3.4 hold and there exists $0<c_{1}<c_{2}<1/2$ such that $c_{1}\leq n_{1}/n\leq c_{2}$. Then for any $f\in\mathcal{F}$ and any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
@@
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq\varepsilon_{\rm BT}+M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
@@
Specifically, when $n_{1}^{*}=O(n_{0}-n_{1})$, $\tilde{n}_{1}=n_{1}^{*}=O(n_{0}-n_{1})$, $n_{0c}=n_{1}^{*}=O(n_{0}-n_{1})$ and $\tilde{n}_{0}=n_{1}^{*}=O(n_{0}-n_{1})$, the result can be written as follows: For any $\alpha>0$ and some constant $C<\infty$, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq\varepsilon_{\rm BT}+C\sqrt{\frac{\log(8/\alpha)}{n_{0}-n_{1}}}.$
@@
Proposition 3.1 establishes an upper bound on the discrepancy between the majority and minority groups. Notably, This upper bound does not directly depend on how accurately the synthetic distribution $\tilde{\mathcal{P}}_{0}$ recovers the true distribution $\mathcal{P}_{0}$. Instead, noting that $\varepsilon_{\rm BT}=2\varepsilon_{h}+2L_{\ell}\varepsilon_{T}$, the discrepancy is governed by three components: the expected loss gap $\varepsilon_{h}$, the transformation approximation error $\varepsilon_{T}$, and the perturbation fluctuation of order $O((n_{0}-n_{1})^{-1/2})$. Since the class probabilities $\pi_{1}=\mathbb{P}(Y=1)$ and $\pi_{0}=\mathbb{P}(Y=0)$ are fixed constants, the effective sample size governing this stochastic fluctuation is of the same order as the total sample size $n$. This result implies that even when the synthetic generator yields a poor approximation to the true minority distribution, the additional error introduced by bias correction remains within the same well-defined bound. In practice, this means that the procedure remains stable and effective under suboptimal synthetic generators, ensuring the reliability of the bias-corrected estimator.
@@
    The loss $\ell_{f}(\bm{x},y)$ is Lipschitz in $\bm{x}$ with constant $L_{\ell}$ uniformly over $y\in\{0,1\}$, i.e., $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2}$ for any $\bm{x}_{1},\bm{x}_{2}\in\mathbb{R}^{d}$ and $y\in\{0,1\}$.
@@
    There exists a constant $C_{3}>0$ such that the following holds for fixed $K$. Let $(\bm{X}_{i})_{i=1}^{n_{1}}$ be an i.i.d. sample drawn from $\mathcal{P}_{1}$. Then for any sample $\bm{x}_{1}$, any $\bm{x}_{2}$ among its $K$ nearest neighbors within the samples, and any $u\in[0,1]$,
    $\ell_{f}(\bm{x}_{1}+u(\bm{x}_{2}-\bm{x}_{1}),1)-\ell_{f}(\bm{x}_{1},1)\geq C_{3}u\|\bm{x}_{2}-\bm{x}_{1}\|_{2}.$
Assumption 3.5 (A1) guarantees that the loss function $\ell_{f}$ is Lipschitz continuous with respect to $\bm{x}$ uniformly over $y\in\{0,1\}$. (A2) ensures that the minority distribution has bounded support and a bounded density function on its support. (A3) holds, for example, if $\ell_{f}(\cdot,1)$ is differentiable on ${\rm supp}(\mathcal{P}_{1})$ and its gradient satisfies $(\bm{x}_{2}-\bm{x}_{1})^{\top}\nabla\ell_{f}(\bm{z},1)\geq C_{3}\|\bm{x}_{2}-\bm{x}_{1}\|_{2}$ for any $\bm{z}$ on the segment between $\bm{x}_{1}$ and $\bm{x}_{2}$.
The following proposition shows that when the minority synthetic samples are generated by SMOTE, the induced bias cannot vanish too quickly. Specifically, there is a lower bound that scales with $(K/n_{1})^{1/d}$, reflecting the discrepancy between $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$.
Proposition 3.2.
Suppose Assumptions 3.1-3.5 hold, and the synthetic minority samples are i.i.d. generated from $\tilde{\mathcal{P}}_{1}$ via SMOTE with parameter $K$. Then there exists a constant $c_{1}>0$, depending on $(d,C_{1},C_{2},C_{3},R)$, such that for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}|\geq c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Proposition 3.2 demonstrates that the bias introduced into the loss function by the synthetic samples is statistically non-negligible. This bias prevents the synthetic augmented loss function (2.2) from serving as a close substitution for the balanced loss function (2.5). This discrepancy between these loss functions leads to a noticeable difference between their corresponding minimizers, which will potentially reduce the performance of the trained classifier.
@@
where $\mathcal{F}=\{f:\mathcal{X}\to(0,1)\}$ represents the prediction function class. In practice, however, one does not have access to the population balanced loss $L^{*}$, but instead minimizes an empirical loss. With synthetic samples generated and the bias correction procedure, we have the empirical minimizer $\widehat{f}^{\rm bc}$ of the bias-corrected loss function (2.9). Next, we investigate the following question: how close is the empirical bias-corrected minimizer $\widehat{f}^{\rm bc}$ to the population balanced minimizer $f^{*}$?
To answer this question, we consider a uniform assumption, which requires that the gap between the bias terms for the minority and majority groups remains controlled. Suppose that Assumption 3.4 is satisfied for all prediction functions $f\in\mathcal{F}$. The basic idea indicates that the bias observed in the minority group can be transferred to the majority group up to a controlled error $\varepsilon_{\rm BT}$, where $\varepsilon_{\rm BT}=2\varepsilon_{h}+2L_{\ell}\varepsilon_{T}$. Intuitively, if $\varepsilon_{\rm BT}$ is small, then correcting for the minority bias using information from the majority group is reliable.
@@
Suppose Assumptions 3.1-3.4 hold and $\tilde{n}_{1}/(n_{0}-n_{1})\to 1$. Then for any $\alpha\in(0,1)$, there exists constant $C$ depends only on $M_{\varepsilon},\pi_{0},\pi_{1}$, such that with probability at least $1-\alpha$,
$L^{*}(\widehat{f}^{\rm bc})-L^{*}(f^{*})\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}+C(2C_{\mathcal{F}}+\sqrt{\log(10/\alpha)})\left\{\frac{1}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right\}.$
@@ 3.2 Bias Correction for Multi-Task Learning
The following theorem establishes nonasymptotic bounds for three types of estimations: the raw MLE $\widehat{\bm{\beta}}_{k}$, the synthetic augmented estimator $\tilde{\bm{\beta}}_{k}$, and the bias-corrected estimator $\tilde{\bm{\beta}}_{k}^{\rm bc}$. It also quantifies how these parameter errors propagate to the estimation of the shared low-rank structure.
@@
Under Assumptions 3.1-3.4 and 3.6, for any $\alpha\in(0,1)$, there exist constants $C_{1},C_{2},C_{3}>0$ such that, with probability at least $1-\alpha$, the following properties hold simultaneously for each source $k$:
${\rm(i)}\quad$ $C_{1}\left(\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}\right)\leq\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}\leq C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\min}(\bm{H}_{k})}\right),$
${\rm(ii)}\quad$ $\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
$\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
${\rm(iii)}\quad$ $\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling},k}\right),$
where $w_{k}=\tilde{n}_{k}/(n_{k}+\tilde{n}_{k1})$ and
$\varepsilon_{{\rm sampling},k}=$ $C_{3}(2C_{\mathcal{F}}+\log(10d/\alpha))\left\{\frac{1}{\sqrt{n_{k1}}}+\frac{1}{\sqrt{n_{k0}}}+\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right\}.$
Furthermore, consider the $d\times K$ true and estimated matrices as $\bm{M}=[\bm{\beta}_{1},\cdots,\bm{\beta}_{K}]$, $\widehat{\bm{M}}^{\rm raw}=[\widehat{\bm{\beta}}_{1}^{\rm raw},\cdots,\widehat{\bm{\beta}}_{K}^{\rm raw}]$, $\widehat{\bm{M}}^{\rm syn}=[\widehat{\bm{\beta}}_{1}^{\rm syn},\cdots,\widehat{\bm{\beta}}_{K}^{\rm syn}]$ and $\widehat{\bm{M}}^{\rm bc}=[\widehat{\bm{\beta}}_{1}^{(\rm bc)},\cdots,\widehat{\bm{\beta}}_{K}^{(\rm bc)}]$. Let $\bm{U}$ be the matrix of leading $r$ left singular vectors of $\bm{M}$ and $D=\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})$ the spectral gap. Then, with the same probability,
${\rm(iv)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm raw},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(v)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm syn},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(vi)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm bc},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2}.$
@@
    The raw MLE $\widehat{\bm{\beta}}_{k}^{\rm raw}$ converges at the standard parametric rate $n_{k}^{-1/2}$.
@@
    The synthetic estimator $\widehat{\bm{\beta}}^{\rm syn}$ inherits an additional bias, which reflects the distributional mismatch between $\mathcal{P}_{k,0}$ and $\tilde{\mathcal{P}}_{k,0}$.
@@
    The bias-corrected estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ removes the bias at the cost of extra sampling fluctuations and the residual class-difference bias $\varepsilon_{\rm BT}$.
@@
the bias correction parameter estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ has smaller errors than the synthetic augmented parameter estimator $\widehat{\bm{\beta}}^{\rm syn}$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed. In addition, when $\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})=O_{p}(n_{k}^{-1/2})$, the lower bound for the synthetic-augmented estimator error $\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}$ in Theorem 3.2 becomes negative and thus can be replaced by zero. Theorem 3.2 shows that the bias correction performs well, especially for “bad” synthetic generators.